• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:05:30 on July 24, 2006

    Entry number 175078

    keyword=Beam Time Accounting
    Hall A

    07/23/06 1600-2400
    Experiment #: CSR
    Shift Leader: AJ Sarty
    Shift Workers: Steffen Strauch, Brandon Craver

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.320.000.000.000.000.000.000.000.32
    Beam Available
    but Not in Use
    0.470.000.000.000.000.000.000.000.47
    Beam Not Available
    or Unacceptable
    0.211.001.001.001.001.001.001.007.21
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    1.000.000.001.001.001.001.001.006.00
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.001.001.000.000.000.000.000.002.00
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.001.001.000.000.000.000.000.002.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: Good production from 16:00-17:00. At 17:00, go to Controlled Access to try & reset HRS-R Q1; at same time, Rad Monitor in BSY tunnel trips and needs to be reset before beam can be delivered. When done with those fixes, no beam is delivered for rest of shift due to vacuum problem in accelerator.

    Data Entry by: AJ Sarty